# UnrealZoo: Enriching Photo-realistic Virtual Worlds for Embodied AI

### v1

@@ Abstract
We introduce UnrealZoo  [^1] Project page: http://unrealzoo.site/, a rich collection of photo-realistic 3D virtual worlds built on Unreal Engine, designed to reflect the complexity and variability of the open worlds. Additionally, we offer a variety of playable entities for embodied AI agents. Based on UnrealCV, we provide a suite of easy-to-use Python APIs and tools for various potential applications, such as data collection, environment augmentation, distributed training, and benchmarking. We optimize the rendering and communication efficiency of UnrealCV to support advanced applications, such as multi-agent interaction. Our experiments benchmark agents in various complex scenes, focusing on visual navigation and tracking, which are fundamental capabilities for embodied visual intelligence. The results yield valuable insights into the advantages of diverse training environments for reinforcement learning (RL) agents and the challenges faced by current embodied vision agents, including those based on RL and large vision-language models (VLMs), in open worlds. These challenges involve latency in closed-loop control in dynamic scenes and reasoning about 3D spatial structures in unstructured terrain.
Figure 1: UnrealZoo enriches photo-realistic virtual worlds by combining diverse scenes and playable entities. It enables training generalizable embodied AI agents for tasks such as navigation, active tracking, and social interactions. Additionally, UnrealZoo facilitates the benchmarking of agents in realistic virtual worlds, helping to identify challenges in open-world deployments.
@@ 1 Introduction
Currently, embodied artificial intelligence (Embodied AI) agents are often homebodies, primarily confined to controlled indoor environments and rarely venturing outside to explore the diversity of the open world. While several simulators (Kolve et al., 2017; Dosovitskiy et al., 2017; Puig et al., 2018; Li et al., 2023; Puig et al., 2024) have advanced the field, they often focus on specific scenarios, such as daily activities in homes or autonomous driving on urban roads. This narrow focus hinders AI agents’ adaptability and generalization in diverse open worlds, e.g., industrial areas, public spaces, and natural landscapes, required by a wide range of real-world applications.
@@
To bridge this gap, there is an increasing demand for simulators that feature a diverse range of open worlds. First, the diversity of the 3D scenes and bodies of agents is crucial to develop spatial intelligence (Davison, 2018), enabling agents to actively perceive, reason, plan, and act when handling numerous tasks in complex 3D worlds with varying dynamics and styles. Second, the complexity of multi-agent interactions is essential for developing social intelligence (Duéñez-Guzmán et al., 2023), such as the theory of mind (Jin et al., 2024), negotiation (Guan et al., 2024), cooperation (Wang et al., 2022), and competition (Zhong et al., 2021), encouraging agents to behave more like humans. Third, virtual worlds that mimic the challenges in open-world scenarios can evaluate agents efficiently and effectively, identifying limitations and preventing hardware losses from real-world deployment failures (Kadian et al., 2020). Ultimately, these features will inspire researchers to explore new challenges previously overlooked in other simulators (Duan et al., 2022), facilitating seamless integration into real-world applications.
@@
In this work, we introduce UnrealZoo, a comprehensive collection of photo-realistic virtual worlds, based on Unreal Engine 222www.unrealengine.com and UnrealCV (Qiu et al., 2017). UnrealZoo features a diverse range of complex open worlds and playable entities to advance research in embodied AI and related domains. This high-quality set includes 100 realistic scenes at varying scales, such as houses, supermarkets, train stations, industrial factories, urban cities, villages, temples, and natural landscapes. Each environment is expertly designed by artists to mimic realistic lighting, textures, and dynamics, closely resembling real-world experiences. Our collection also includes diverse playable entities, including humans, animals, robots, drones, motorbikes, and cars. This diversity enables researchers to investigate the generalization of agents across different embodiments or build complex 3D social worlds with numerous heterogeneous agents. To enhance usability, we further optimize UnrealCV and offer a suite of easy-to-use Python APIs and tools (UnrealCV+), including environment augmentation, demonstration collection, and distributed training/testing. These tools allow for customization and extension of the environments to meet various needs in future applications, ensuring UnrealZoo remains adaptable as the embodied AI agents evolve.
@@
We conduct experiments to demonstrate the applicability of UnrealZoo for embodied AI. First, we benchmark frames per second (FPS) across various commands, highlighting the significant improvement in image rendering and multi-agent interactions with the UnrealCV+ API. We use embodied visual navigation (Zhu et al., 2017) and tracking (Luo et al., 2018; Zhong et al., 2021) as two example tasks to benchmark embodied vision agents in complex dynamic environments with moving objects and unstructured maps. We also introduce a set of simple yet effective baseline methods for developing embodied vision agents, including distributed online reinforcement learning algorithms (Mnih et al., 2016), offline reinforcement learning algorithms (Kumar et al., 2020), and a reasoning framework for large vision-language models (VLMs). Our evaluations across different settings emphasize the importance of diverse training environments for enhancing agent generalization and robustness, the necessity of low latency in closed-loop control to handle dynamic factors, and the potential of reinforcement learning for training agents to navigate complex scenes.
@@
Our contributions can be summarized in the following: 1) We build UnrealZoo, a collection of 100 high-quality photo-realistic scenes and a set of playable entities with diverse features, covering the most challenging to embodied AI agents in open worlds. 2) We optimize the communication efficiency of UnrealCV APIs and provide easy-to-use Gym interfaces with a toolkit for diverse requirements. 3) We conduct experiments to demonstrate the usability of UnrealZoo, showing the importance of the diversity of the environments to the embodied agents, and analyzing the limitations of the current RL-based and VLM-based agents in the open worlds.
@@ 2 Related Works
Realistic Simulators for Embodied AI. Realistic simulators are extensively utilized in embodied artificial intelligence due to their appealing benefits, including high-quality rendering, cost-effective ground truth generation, low-cost interaction, and environmental controllability. They are crucial for training and testing AI agents to handle increasingly complex tasks. Notable realistic 3D simulators have been created for specific applications, such as indoor navigation (Kolve et al., 2017; Puig et al., 2018; Xia et al., 2018; Wu et al., 2018), robot manipulation (Yu et al., 2020; Ehsani et al., 2021; Chen et al., 2024), and autonomous driving (Gaidon et al., 2016; Shah et al., 2018; Dosovitskiy et al., 2017). Recent advances in computer graphics have spurred interest in developing general-purpose virtual worlds with photo-realistic rendering, allowing agents to collect high-fidelity data and learn skills applicable across various tasks and scenes. ThreeDWorlds (TDW) (Gan et al., 2021) and LEGENT (Cheng et al., 2024) are notable simulators that offer photo-realistic, multi-modal platforms, based on Unity, for interactive physical simulation. However, their built-in scenes and playable entities are somewhat limited. Additionally, the performance of the simulator decreases significantly in large outdoor environments, a typical weakness of Unity. V-IRL (Yang et al., 2024) is a recent approach that leverages Google Maps’ API to simulate agents with real-world street view images, significantly reducing the gap between virtual and real-world settings. However, since V-IRL is inherently composed of static images, it lacks the capability to simulate the dynamics of the physical worlds for agent-object interactions. Recently, the community has also begun to explore dynamic environments with social interactions and unexpected events. However, existing solutions like Habitat 3.0 (Puig et al., 2024) focus on a limited number of agent interactions in indoor scenes, while HAZARD (Zhou et al., 2024b) addresses only single-agent simulations in dynamic scenarios like fires, floods, and winds. In contrast, UnrealZoo offers a comprehensive collection of scenes that feature various scenes at different scales, situations, eras, and cultural backgrounds with diverse playable entities for embodied AI. With advancements in Unreal Engine and optimized UnrealCV, our environment achieves real-time performance in large-scale scenes with multiple agents (around 10) and photo-realistic rendering. A comprehensive comparison across the related photo-realistic simulators is shown in Table 1.
Table 1: The comparison with related photo-realistic virtual worlds for embodied AI. Unstr. Terr. indicates the presence of unstructured terrain. Nav. Sys. specifies whether the agent in the environment includes an autonomous navigation system. In Appendix A, we compare the visual realism across different engines in Figure 6 and list the descriptions of the symbols in Table 6.
@@
Embodied Vision Agents. Embodied vision agents, which perceive and interact with their environments through vision, are a key focus in artificial intelligence research. These agents perform tasks like navigation (Zhu et al., 2017; Gupta et al., 2017; Yokoyama et al., 2024; Long et al., 2024), active object tracking (Luo et al., 2018; Zhong et al., 2019; 2021; 2023; 2024), and other interactive tasks (Chaplot et al., 2020; Weihs et al., 2021; Ci et al., 2023; Wang et al., 2023), mimicking human behavior. Their development involves various methods, including state representation learning (Yadav et al., 2023; Yuan et al., 2022; Gadre et al., 2022; Yang et al., 2023), reinforcement learning (RL) (Schulman et al., 2017; Xu et al., 2024; Ma et al., 2024), and large vision-language models (VLMs) (Zhang et al., 2024; Zhou et al., 2024a). Despite significant progress, challenges remain. RL methods often require extensive trial-and-error interactions and computational resources for training, and they usually struggle to generalize to new environments. Conversely, VLM-based methods excel at interpreting language instructions and images but may lack the fine-grained control and adaptability necessary for real-time interactions. The computational demands and time needed for inference with such large models are critical, especially in dynamic scenes. Moreover, previous simulators mainly focus on indoor rooms or urban roads, which mask the potential challenges to the embodied agents when deploying in open worlds, e.g., unstructured terrain, dynamic changing factors, inference costs of the perception-control loop, and social interactions with other agents. Therefore, it is required to benchmark agents in large-scale, photo-realistic open worlds, taking into account various real-world challenges in the virtual worlds. In this work, we collect a subset of environments from UnrealZoo and benchmark embodied visual navigation and tracking agents, to emphasize the weakness of the existing RL-based and VLM-based methods.
@@ 3 UnrealZoo
UnrealZoo is a collection of photo-realistic, interactive open-world environments with diverse embodied characters, built on Unreal Engine and UnrealCV (Qiu et al., 2017). The environments are sourced from the Unreal Engine Marketplace  [^3] https://www.unrealengine.com/marketplace, which shares high-quality content from artists, and were accumulated over two years at a cost exceeding $10,000$. UnrealZoo features a diverse array of scenes with varying sizes and styles. Among them, the largest scene, i.e., Medieval Nature Environment, covers more than $16km^{2}$ areas. The environments also include a wide range of embodiment, such as human avatars, vehicles, drones, animals, and virtual cameras, all of which can interact with the environment and equipped with ego-centric sensing systems. We offer easy-to-use Python APIs based on UnrealCV to facilitate interaction between Python programs and the game engine. Note that UnrealCV is optimized for rendering and communication, particularly in large-scale and multi-agent scenarios, namely UnrealCV+. Additionally, we provide OpenAI Gym interfaces to standardize agent-environment interactions. The gym-like interface also contains a set of toolkits, e.g., environment augmentation, population control, time dilation, and JSON-style task configurations to help the user customize the environments for various tasks with minimal effort.
@@ 3.1 Scene Collection
UnrealCV Zoo contains 100 scenes based on Unreal Engine 4 and 5. We select the scene based on the public reviews in the marketplace and the difference between the collected scenes, aiming at covering a wide range of styles from ancient to fictional, ensuring diversity. We provide an overview of the environments in the scene gallery.
Figure 2: The detailed architecture of UnrealZoo. The Gray box indicates the UE binary, collecting the scenes and playable entities. The UnrealCV+ Server is built in the binary as a plugin. We have bolded the names of the optimized or new modules in UnrealCV+ Server and Client. We provide OpenAI Gym Interfaces for agent-environment interaction, which has been widely used in the community. Our gym interface supports customizing the task in a configuration file and contains a toolkit with a set of gym wrappers for environment augmentation, population control, etc.
We have tagged the collected scenes with several feature labels allowing researchers to select appropriate scenes for testing or training based on the tags associated with each scene. Our tags cover the following aspects:
@@
  • Style: The scenes may also represent diverse styles that reflect various cultures and eras, such as Asian Temple, Western Church, and Middle Eastern Street. Cultural labels include Western, Asian, and Middle Eastern, while era labels encompass Medieval, Modern, and Science Fiction. Identifying the styles will help us build a new data set to benchmark how social agents adapt to different backgrounds.
After categorizing the scenes, we integrate UnrealCV+ (Refer to Section 3.3) into the UE project and add the controllable player assets (Refer to Section 3.2) to each scene. Due to licensing restrictions, content purchased from the marketplace cannot be open-source, so we package the projects into an executable binary for sharing with the community. These executable binaries will be compatible with various operating systems, including Windows, Linux, and macOS, allowing users to download and run them via the Python interface without needing any knowledge of Unreal Engine, which is primarily built on C++ and Blueprint.
@@ 3.2 Playable Entities
Figure 3: First-person (Top) and third-person (Bottom) view images of different entities in different scenes. Note that camera parameters can be reconfigured by UnrealCV APIs.
@@
The locomotion system is built on Smart Locomotion, a well-designed and smooth locomotion system. It contains a number of high-quality animations that enable the agent to interact with the scene, such as opening and closing doors, crouching under obstacles, jumping over obstacles, climbing onto a platform, and simulating injury or death. With the locomotion system, we can explore the agent’s ability to reason, plan, and interact in large-scale complex 3D scenes in advance, ignoring learning skills for low-level action control that requires high-fidelity physical simulation.
The navigation system is built on NavMesh allowing agents to autonomously navigate with the built-in AI controller. This includes path-finding and obstacle-avoidance capabilities, ensuring smooth and realistic movement throughout diverse terrains and structures. For urban-style maps, we segment the roads to distinguish between pedestrian and vehicle pathways. When agents use the navigation system for autonomous control, they will navigate the shortest path based on the priority of the different areas. For example, pedestrians and animals will prioritize walking on sidewalks, while vehicles and motorcycles will prioritize driving on roadways. An example of the navigation area is shown in Figure 10.
@@ 3.3 Programming Interface
We provide UnrealCV+ as the basic application programming interface (API) on Python to capture data and control the entities and scenes, and provide an OpenAI Gym interface for general agent-environment interactions. The architectures of the programming interfaces are shown in Figure 2.
UnrealCV+ is our improved version of the UnrealCV (Qiu et al., 2017) for high-throughput interactions. As the original version of UrnealCV primarily focuses on generating synthesis data for computer vision, the frame rates per second (FPS) are not optimized for real-time interactions. We optimize the rendering pipelines in the UnrealCV server and the communication protocols between the server and the client to improve the FPS. Specifically, we enable parallel processing while rendering object masks and depth images, which can significantly improve the FPS in large-scale scenes. For multi-agent interactions, we further introduce the batch commands protocol. In this protocol, the client can simultaneously send a batch of commands to the server, processing all the received commands and returning a batch of results. In this way, we can reduce the time spent on server-client communication. Since reinforcement learning requires extensive trial-and-error interactions for training, often running multiple environments on a computer, we introduce Inter-process communication (IPC) sockets instead of TCP sockets to improve the stability of the server-client communication under high loads. We benchmark the FPS performance in Table 2. To enhance user-friendliness, we have developed high-level Python APIs that are built upon the command systems of UnrealCV. These APIs encapsulate all the request commands and their corresponding data decoders into a callable Python function. This approach significantly simplifies the process for beginners, allowing them to interact with and customize the environment using UnrealCV+.
Table 2: Comparison of FPS in Unreal Engine 4.27 with UnrealCV and UnrealCV+. The reported result is an average performance across 6 typical environments.
@@
Gym Interface is used to define the interactive tasks and standardize the agent-environment interaction, following Gym-UnrealCV. Even though there are a lot of tasks for agents, they usually share common interaction protocols, i.e., the agent gets observations from the environment and returns actions. The main difference across different tasks usually is the reward functions, the modality of the observation, and the available actions. Hence, we define the basic interaction functions for general usage and list the task-specific configurations, e.g., scene name, and reward function, in a JSON File, as shown in Figure 12. In this way, when adding new UE scenes, the users only need to set the parameters in the JSON files. Moreover, we contain a toolkit with a set of gym wrappers for training and testing the agents, such as environment augmentation that has been in previous work for training generalizable agents (Luo et al., 2018; 2020), population control to adjust the number of agents in the scene, and time dilation to adjust the control frequency in dynamic scenes. In Section 4.3, we demonstrate an example usage of the toolkit to analyze the robustness of social tracking agents to the population of crowds and the impact of the control frequency in such dynamic scenes. We also provide a launch tool to enable the user to run multiple environments with specific GPU IDs within a computer, which is useful for distributed online reinforcement learning (Ci et al., 2023).

### v2

@@ Abstract
We introduce UnrealZoo, a collection of over 100 photo-realistic 3D virtual worlds built on Unreal Engine, designed to reflect the complexity and variability of open-world environments. We also provide a rich variety of playable entities, including humans, animals, robots, and vehicles for embodied AI research. We extend UnrealCV with optimized APIs and tools for data collection, environment augmentation, distributed training, and benchmarking. These improvements achieve significant improvements in the efficiency of rendering and communication, enabling advanced applications such as multi-agent interactions. Our experimental evaluation across visual navigation and tracking tasks reveals two key insights: 1) environmental diversity provides substantial benefits for developing generalizable reinforcement learning (RL) agents, and 2) current embodied agents face persistent challenges in open-world scenarios, including navigation in unstructured terrain, adaptation to unseen morphologies, and managing latency in the close-loop control systems for interacting in highly dynamic objects. UnrealZoo thus serves as both a comprehensive testing ground and a pathway toward developing more capable embodied AI systems for real-world deployment.
Figure 1: UnrealZoo enriches photo-realistic virtual worlds with diverse scenes and playable entities, enabling training of generalizable embodied AI agents for navigation, tracking, and social interactions.
@@ 1 Introduction
Currently, embodied artificial intelligence (Embodied AI) agents are often homebodies, primarily confined to controlled indoor environments and rarely venturing outside to explore the diversity of the open world. While several simulators [20, 8, 28, 22, 29] have advanced the field, they often focus on specific scenarios, such as daily activities in homes or autonomous driving on urban roads. This narrow focus hinders AI agents’ adaptability and generalization in open worlds, such as factories, public areas, and landscapes, which are required by several real-world applications.
@@
To bridge this gap, there is a growing demand for simulators that feature diverse open-world environments. First, the diversity of 3D scenes and morphologies is fundamental to developing spatial intelligence  [7]. This diversity enables agents to actively perceive, reason, plan, and act while tackling various tasks in complex 3D worlds. Second, the complexity of multi-agent interactions is essential for developing social intelligence [10], such as the theory of mind [17], negotiation [15], cooperation [34], and competition [47], encouraging agents to behave more like humans. Third, virtual worlds that mimic the challenges in open-world scenarios can evaluate agents efficiently and effectively, identifying limitations and preventing hardware losses from real-world deployment failures [18]. Ultimately, these features will inspire researchers to explore new challenges previously overlooked in other simulators [9], facilitating seamless integration into real-world applications.
@@
In this work, we introduce UnrealZoo, a comprehensive collection of photo-realistic virtual worlds, based on Unreal Engine and UnrealCV [30]. UnrealZoo features a diverse range of complex open worlds and playable entities to advance research in embodied AI and related domains. This high-quality set includes 100 realistic scenes at varying scales, such as houses, supermarkets, train stations, industrial factories, urban cities, villages, temples, and natural landscapes. Each environment is expertly designed by artists to mimic realistic lighting, textures, and dynamics, closely resembling real-world experiences. Our collection also includes diverse playable entities, including humans, animals, robots, drones, motorbikes, and cars. This diversity enables researchers to investigate the generalization of agents across different embodiments or build complex 3D social worlds with numerous heterogeneous agents. To enhance usability, we further optimize UnrealCV and offer a suite of easy-to-use Python APIs and tools (UnrealCV+), including environment augmentation, demonstration collection, and distributed training/testing. These tools allow for customization and extension of the environments to meet various needs in future applications, ensuring UnrealZoo remains adaptable as the embodied AI agents evolve.
@@
We conduct experiments to demonstrate the applicability of UnrealZoo for embodied AI. First, we benchmark frames per second (FPS) across various commands, highlighting the significant improvement in image rendering and multi-agent interactions with the UnrealCV+ API. We use embodied visual navigation [52] and tracking [24, 47] as two example tasks to benchmark embodied vision agents in complex dynamic environments with moving objects and unstructured maps. We also introduce a set of simple yet effective baseline methods for developing embodied vision agents, including distributed online reinforcement learning algorithms [27], offline reinforcement learning algorithms [21], and a reasoning framework for large vision-language models (VLMs). Our evaluations across different settings emphasize the importance of diverse training environments for enhancing agent generalization and robustness, the necessity of low latency in closed-loop control to handle dynamic factors, and the potential of reinforcement learning for training agents to navigate complex scenes.
@@
Our contributions can be summarized as follows: 1) We build UnrealZoo, a collection of 100 high-quality photo-realistic scenes along with a set of playable entities. These cover the most challenging aspects for embodied AI agents in open-world environments. 2) We optimize the communication efficiency of UnrealCV APIs and provide easy-to-use Gym interfaces with a toolkit to meet diverse requirements. 3) We conduct experiments to demonstrate the usability of UnrealZoo, showing the importance of the diversity of the environments to the embodied agents, and analyzing the limitations of the current RL-based and VLM-based agents in the open worlds.
Table 1: The comparison with related photo-realistic virtual worlds for embodied AI. Nav. Sys. specifies whether the agent in the environment includes an autonomous navigation system. We compare the visual realism across different engines and list the descriptions of the symbols in supplementary materials.
@@ 2 Related Works
Realistic Simulators for Embodied AI. Realistic simulators are widely adopted in embodied AI research due to their high-quality rendering, cost-effective ground truth generation, and environmental controllability. Specialized 3D simulators have been developed for indoor navigation [20, 28, 37, 36], robot manipulation [43, 11, 4], and autonomous driving [13, 32, 8]. Recent advances in computer graphics have enabled general-purpose virtual worlds with photo-realistic rendering. ThreeDWorlds (TDW) [14] and LEGENT [5] offer multi-modal platforms based on Unity, but their scenes and playable entities remain limited, with performance degrading in large outdoor environments. V-IRL [41] leverages Google Maps’ API to simulate agents with real-world street view images, yet lacks physical interaction capabilities due to its static image composition. While the community has begun exploring dynamic environments with social interactions, existing solutions like Habitat 3.0 [29] focus on limited agent interactions in indoor scenes, and HAZARD [51] addresses only single-agent simulations in scenarios like fires and floods. In contrast, UnrealZoo offers a comprehensive collection of diverse scenes across different scales, situations, eras, and cultural backgrounds with various playable entities. Leveraging Unreal Engine advancements and optimized UnrealCV, our environment achieves real-time performance in large-scale scenes with multiple agents and photo-realistic rendering. A detailed comparison of photo-realistic simulators is presented in Table 1.
@@
Embodied Vision Agents. This represents a key frontier in AI research, performing tasks with visual observation, like navigation [52, 16, 42, 23], active object tracking [24, 46, 47, 48, 49], and other interactive tasks [2, 35, 6, 33]. Their development involves state representation learning [39, 44, 12, 40], reinforcement learning (RL) [31, 38, 26], and large vision-language models (VLMs) [45, 50], such as GPT-4o and SpatialVLM  [3]. Despite advances, significant challenges persist. RL methods require extensive trial-and-error interactions and computational resources, often struggling with generalization. VLM-based methods excel at interpreting instructions and images but may lack fine-grained control and adaptability for real-time interactions. Importantly, previous simulators primarily focus on indoor rooms or urban roads, overlooking critical challenges faced by embodied vision agents in open worlds: unstructured terrain, dynamic factors, perception-control loop costs, and multi-agent social interactions. Therefore, benchmarking agents in expansive, photo-realistic virtual environments is essential to address real-world challenges. In this work, we collect a subset of environments from UnrealZoo and benchmark embodied visual navigation and tracking agents, revealing limitations in current methods that were previously masked by constrained simulation environments.
@@ 3 UnrealZoo
UnrealZoo is a collection of photo-realistic, interactive open-world environments with diverse embodied characters, built on Unreal Engine and UnrealCV [30]. The environments are sourced from the Unreal Engine Marketplace  [^1] https://www.unrealengine.com/marketplace, which shares high-quality content from artists, and were accumulated over two years at a cost exceeding $10,000$. UnrealZoo features a diverse array of scenes with varying sizes and styles. Among them, the largest scene, i.e., Medieval Nature Environment, covers more than $16km^{2}$ areas. The environments also include a wide range of embodiment, such as human avatars, vehicles, drones, animals, and virtual cameras, all of which can interact with the environment and are equipped with ego-centric sensing systems. We offer easy-to-use Python APIs based on UnrealCV to facilitate interaction between Python programs and the game engine. Note that UnrealCV is optimized for rendering and communication, particularly in large-scale and multi-agent scenarios, namely UnrealCV+. Additionally, we provide OpenAI Gym interfaces to standardize agent-environment interactions. The gym-like interface also contains a set of toolkits, e.g., environment augmentation, population control, time dilation, and JSON-style task configurations to help the user customize the environments for various tasks with minimal effort.
@@ 3.1 Scene Collection
UnrealCV Zoo contains 100 scenes based on Unreal Engine 4 and 5. We select the scene based on the public reviews in the marketplace and the difference between the collected scenes, aiming at covering a wide range of styles from ancient to fictional, ensuring diversity. We provide an overview of the environments in the scene gallery, and the detailed statistical distribution is shown in Figure  2.
We have tagged the collected scenes with several feature labels, allowing researchers to select appropriate scenes for testing or training based on the tags associated with each scene. Our tags cover the following aspects:
@@
  • Style: The scenesen compass diverse cultural and historical styles, including Asian Temple, Western Church, and Middle Eastern Street. Cultural labels include Western, Asian, and Middle Eastern, while era labels encompass Medieval, Modern, and Science Fiction. Identifying the styles will help us build a new data set to benchmark how social agents adapt to different backgrounds.
After categorizing the scenes, we integrate UnrealCV+ (Refer to Section 4) into the UE project and add the controllable player assets (Refer to Section 3.2) to each scene. Due to licensing restrictions, content purchased from the marketplace cannot be open-source, so we package the projects into an executable binary for sharing with the community. These executable binaries will be compatible with various operating systems, including Windows, Linux, and macOS, allowing users to download and run them via the Python interface without needing any knowledge of Unreal Engine, which is primarily built on C++ and Blueprint.
@@ 3.2 Playable Entities
Figure 3: First-person and third-person view images of different entities in different scenes. Note that camera parameters can be reconfigured by UnrealCV APIs.
@@
The locomotion system is built on Smart Locomotion, an animation framework providing natural movement and environmental interactions. Interactions are animation-based rather than physics-based, featuring binary state changes (e.g., opening doors without physical doorknob manipulation) and positional animations (crouching, jumping, climbing), similar to VirtualHome [28]. Interactive objects maintain consistent behavior across all environments and can be spawned at arbitrary locations with customizable visual properties via the API. This approach facilitates research on high-level reasoning and planning without the need for extra implementation of primitive motion control.
The navigation system is built on NavMesh (a spatial partitioning system that defines walkable surfaces), allowing agents to autonomously navigate with the built-in AI controller. This includes path-finding and obstacle-avoidance capabilities, ensuring smooth and realistic movement throughout diverse terrains and structures. For urban-style maps, we segment the roads to distinguish between pedestrian and vehicle pathways. When agents use the navigation system for autonomous control, they will navigate the shortest path based on the priority of the different areas. For example, pedestrians and animals will prioritize walking on sidewalks, while vehicles and motorcycles will prioritize driving on roadways. An example of the navigation area is shown in Supplementary Materials.
@@ 3.3 Programming Interface
Figure 4: The detailed architecture of UnrealZoo. The Gray box indicates the UE binary, containing the scenes, playable entities, and UnrealCV+ server. We have bolded the names of the optimized or new modules in UnrealCV+ Server and Client. We provide OpenAI Gym Interfaces for agent-environment interaction. Our gym interface supports customizing the task in a configuration file and contains a toolkit with a set of gym wrappers for environment augmentation, population control, and etc.
We provide UnrealCV+ as the fundamental interface on Python to capture data and control the entities and scenes, and provide an OpenAI Gym interface  [^2] https://github.com/UnrealZoo/unrealzoo-gym for general agent-environment interactions. The architectures of the programming interfaces are shown in Figure 4.
Table 2: Comparison of FPS in Unreal Engine 4.27 with UnrealCV and UnrealCV+. Results represent average performance across 6 typical environments ranging from small-scale scenes ($2,440m^{2}$) to expansive landscapes ($16km^{2}$). Benchmarks were conducted on a system with an Nvidia RTX 4090 GPU, Intel i7-14700k CPU, and Windows OS. Performance metrics were collected by executing each Python function 1,000 times and measuring execution duration. Multi-agent interaction testing utilized the gym interface, with all rendered images at 640×480 resolution. See the supplementary material for the detailed results.
@@
UnrealCV+ is our improved version of the UnrealCV [30] for high-throughput interactions. As the original version of UrnealCV primarily focuses on generating synthesis data for computer vision, the frame rates per second (FPS) are not optimized for real-time interactions. We optimize the rendering pipelines in the UnrealCV server and the communication protocols between the server and the client to improve the FPS. Specifically, we enable parallel processing while rendering object masks and depth images, which can significantly improve the FPS in large-scale scenes. For multi-agent interactions, we further introduce the batch commands protocol. In this protocol, the client can simultaneously send a batch of commands to the server, processing all the received commands and returning a batch of results. In this way, we can reduce the time spent on server-client communication. Since reinforcement learning requires extensive trial-and-error interactions, often running multiple environments on a computer, we introduce Inter-process communication (IPC) sockets instead of TCP sockets to improve the stability of the server-client communication under high loads. We benchmark the FPS performance in Table 2. To enhance user-friendliness, we have developed high-level Python APIs that are built upon the command systems of UnrealCV. These APIs encapsulate all the request commands and their corresponding data decoders into a callable Python function. This approach significantly simplifies the process for beginners, allowing them to interact with and customize the environment using UnrealCV+.
Gym Interface is used to define the interactive tasks and standardize the agent-environment interaction, following Gym-UnrealCV. Even though there are a lot of tasks for agents, they usually share common interaction protocols, i.e., the agent gets observations from the environment and returns actions. The main difference across different tasks is usually the reward functions, the modality of the observation, and the available actions. Hence, we define the basic interaction functions for general usage and list the task-specific configurations, e.g., scene name, and reward function, in a JSON File. In this way, when adding new UE scenes, the users only need to set the parameters in the JSON files. Moreover, we provide a toolkit with a set of gym wrappers for training and testing the agents, such as environment augmentation that has been in previous work for training generalizable agents [24, 25], population control to adjust the number of agents in the scene, and time dilation to adjust the control frequency in dynamic scenes. In Section 4.3, we demonstrate an example usage of the toolkit to analyze the robustness of social tracking agents to the population of crowds and the impact of the control frequency in such dynamic scenes. We also provide a launch tool to enable the user to run multiple environments with specific GPU IDs within a computer, which is useful for distributed online reinforcement learning.